OpenStudy (anonymous):

Find the exact value of cos-1the quantity square root of three divided by two.

OpenStudy (anonymous):

\[\cos^{-1} (\sqrt{3}/2)\]

Nnesha (nnesha):

we take inverse of ratio to find an angle so you need an angle when cos theta = sqrt{3}/2 1st) you can use unit circle 2nd) calculator 3rd) familiar with 30-6-90 /45-45-90 triangle theorem

Nnesha (nnesha):

btw in the order pair (x,y) cos = x sin = y

OpenStudy (anonymous):

so what do i do next

Nnesha (nnesha):

1st) you can use unit circle 2nd) calculator 3rd) familiar with 30-6-90 /45-45-90 triangle theorem 3 ways to find an answer pick one :=)

Nnesha (nnesha):

it's 30-60-90 ****

OpenStudy (anonymous):

unit circle i guess

Nnesha (nnesha):

alright  at what angle cos = sqrt{3}/2

OpenStudy (anonymous):

30 and 60

OpenStudy (anonymous):

no just 30

Nnesha (nnesha):

there should be two angle when cos =sqrt{3}/2 remember it's reflected over they axis 

Nnesha (nnesha):

the*

OpenStudy (anonymous):

i thought you only needed one but okay then 330

Nnesha (nnesha):

well there are two you should write both as an ansaawer just in case if they are looking for more than one value hmm

OpenStudy (anonymous):

okay so whats the next step?

Nnesha (nnesha):

that's it

Nnesha (nnesha):

is there any statement about degree or radian ? i mean answer should be in degree or radian ?

OpenStudy (anonymous):

but its the arc cosine

OpenStudy (anonymous):

its a radian

Nnesha (nnesha):

right when we need an angle we take invs of ratio so that's why it's arccos

Nnesha (nnesha):

ohh okay formula to convert degree to radian \[\huge\rm rad=\frac{ \pi}{ 180} \times degree \]

OpenStudy (anonymous):

well which angle do i use

Nnesha (nnesha):

both angles 30 and 330

OpenStudy (anonymous):

cant i just use the radian it has on the unit circle next to it

Nnesha (nnesha):

yes you can

Nnesha (nnesha):

you will get the same answer

OpenStudy (anonymous):

but both are an answer choice

OpenStudy (anonymous):

which one is right

Nnesha (nnesha):

ohh so you can pick only one ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

i'll pick 30 bec that's the angle where it starts

Nnesha (nnesha):

330 is areflection

OpenStudy (anonymous):

so pi/6?

Nnesha (nnesha):

looks good

OpenStudy (anonymous):

thank you

Nnesha (nnesha):

yw